OpenStudy (anonymous):

I vaguely remember my teacher saying that any triangle that is inscribed in a semi circle is always a right triangle. is that true? because i kinda proved him wrong, or i heard something wrong... is there a specific rule for that?

OpenStudy (anonymous):

yes, its correct

OpenStudy (anonymous):



OpenStudy (anonymous):

 2A+C =180 2B+D=180 2A+2B +C+D=360 but C+D = 180, so: 2A+2B+180=360 2A+2B = 180 A+B = 90 which prooves the staement